OpenStudy (anonymous):

The Partial Derivative Test find the critical values for f(x,y)=6xy^2 -2x^3 - 3y^4 test and classify the nature of the extrema

OpenStudy (serenity74):

Calculus is life lol you don't question you just do it XD JK.

OpenStudy (zarkon):

what have you tried?

OpenStudy (anonymous):

fx=6y^2-6x^2 fy=12xy-12y^3

OpenStudy (anonymous):

i set fx equal to 0 and got x=y

OpenStudy (zarkon):

can you first spot the "obvious" solution

OpenStudy (anonymous):

no

OpenStudy (zarkon):

notice x=0 and y=0 is a solution

OpenStudy (zarkon):

then \[f_y=12xy-12y^3 =0\] \[y(12x-12y^2)=0\] so \[12x-12y^2=0\]

OpenStudy (zarkon):

\[f_x=6y^2-6x^2=0 \] so \[12y^2-12x^2=0 \]

OpenStudy (zarkon):

add \[12x-12y^2=0\] and \[12y^2-12x^2=0\] then solve for x

OpenStudy (anonymous):

what do you mean add?

OpenStudy (anonymous):

if we know x=0 and y=0 then (0,0)

OpenStudy (anonymous):

shouldnt we use the determinant fxx*fyy-(fx)^2 to test its extrema?